Leno Company manufactures toasters. For the first 8 months of 2014, the company reported the following operating results while operating at 75% of plant capacity Sales (350,500 units) $4,372,000 Cost of goods sold 2,599,100 Gross profit 1,772,900 839,200 $933,700 Net income Cost of goods sold was 68% variable and 32% fixed; operating expenses were 73% variable and 27% fixed. In September, Leno Company receives a special order for 17,400 toasters at $8.05 each from Centro Company of Ciudad Juarez. Acceptance of the order would result in an additional $3,040 of shipping costs but no increase in fixed operating expenses. Reject Order Accept Order Net Income Increase (Decrease) Cost of goods sold Operating expenses Net income SHOW SOLUTION SHOW ANSWER Should Leno Company accept the special order?
Answer #1

Fixed cost of goods sold = 2,599,100 x 32%

= $831,712

Variable cost of goods sold for 350,500 units = 2,599,100 - 831,712

= $1,767,388

Hence, variable cost of goods sold per unit = 1,767,388/350,500

= $5.04

Cost of goods sold when special order is accepted = (1,767,388 + 17,400 x 5.04) + 831,712

= $2,686,839

Fixed operating expenses = 839,200 x 27%

= $226,584

Variable operating expenses for 350,500 units = 839,200 - 226,584

= $612,616

Hence, variable operating expense per unit = 612,616/350,500

= $1.75

Additional shipping cost of $3,040 will be incurred if order is accepted.

Hence, operating expenses if order is accepted = (612,616 + 17,400 x 1.75 + 3,040) + 226,584

= $872,690

Sales revenue when order is accepted = 4,372,000 + 17,400 x 8.05

= $4,512,070

Reject order Accept order Net income increase (decrease)
Revenues 4,372,000 4,512,070 140,070
Cost of goods sold 2,599,100 2,686,839 (87,739)
Operating expenses 839,200 872,690 (33,490)
Net income 933,700 952,541 18,841

If special order is accepted, net income will increase by $18,841. Hence special order must be accepted.

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.